John is three times his son's age. If the difference between their ages is 30, how old is the son.
I think the problem is set up like 3n-n=30 If this is right, how do I solve it?

Answer by stanbon(75887)

John is three times his son's age.
If the difference between their ages is 30, how old is the son.
----
Equations:
j = 3s
j-s = 30
-------------
Substitute for "j" and solve for "x":
3s-s = 30
2s = 30
s = 15 (The son's age now)
---
Solve for "j":
j = 3s
j = 3*15 = 45 yrs old (John's age now)
===================================
Cheers,
Stan H.
========================